Raquietball. The volume of rubber(in cubic centimeters) in a hollow rubber ball used the racquetball is given by V = 4/3 pie R to the third power - 4/3 pie r to the third power. where the inside radius is r centimeters and the outside radius ir R centimeters.
a) Rewrite the formula by factoring the right-hand side compeletely.
b) The accompanying graph shows the relationship between r and V when R=3. Use the graph to estimate the value of r for which V=100 cm cubed.

Answer by stanbon(48558)

Raquietball. The volume of rubber(in cubic centimeters) in a hollow rubber ball used the racquetball is given by V = 4/3 pie R to the third power - 4/3 pie r to the third power. where the inside radius is r centimeters and the outside radius ir R centimeters.
a) Rewrite the formula by factoring the right-hand side compeletely.
b) The accompanying graph shows the relationship between r and V when R=3. Use the graph to estimate the value of r for which V=100 cm cubed.
Volume of the rubber = (4/3)(pi)(R^3-r^3)
100 = (4/3)(pi)(27-r^3)
300/(4pi)=27-r^3
23.87=27-r^3
r^3=3.12675...
r=1.462 cm
Cheers,
stan H.